### v1

# Intensified optical camera with Timepix4 readout

Abstract
We report the first characterization results of an optical time-stamping camera based on the Timepix4 chip coupled to a fully depleted optical silicon sensor and fast image intensifier, enabling sub-nanosecond scale, time-resolved imaging for single photons. The system achieves an RMS time resolution of 0.27 ns in direct detection mode without the intensifier and from 0.55 to 1.4 ns in the single-photon regime with intensifier for different amplitude-based signal selections. This shows that Timepix4 provides a significant improvement over previous Timepix3-based cameras in terms of timing precision, pixel count, and data throughput. We analyze key factors that affect performance, including sensor bias and time-walk, and demonstrate effective correction methods to recover high temporal accuracy. The camera’s temporal resolution, event-driven readout and high rate capability make it a scalable platform for a wide range of applications, including quantum optics, ultrafast imaging, and time-correlated photon counting experiments.
keywords:
Timepix optical camera, Timepix3CAM, timewalk calibration, Timepix3, Timepix41 Introduction
The ability to detect, count, and time-stamp individual optical photons is becoming increasingly important across a wide range of imaging applications. This approach enables the capture of the complete spatio-temporal and, in principle, also spectral information carried by each detected photon, providing deeper insight into the underlying physical processes. Photon-counting is already well established in X-ray imaging, where sufficient photon energy allows for direct detection and time-resolved, energy-resolved measurements. High-rate readout electronics further enable rapid accumulation of statistics, making this modality highly efficient for quantitative imaging.
In the optical domain, the development of time-stamping cameras — particularly those based on the Timepix family readout chips — has enabled similar capabilities for visible and near-infrared photons [1, 2]. These systems offer nanosecond-scale timing resolution, opening new avenues in areas such as ion imaging [3, 4, 5, 6], optical readout of time-projection chambers (TPC) [7], fluorescence lifetime imaging [8, 9, 10], neutron detection [11, 12, 13, 14] and quantum sciences [15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. Timepix-based single photon sensitive cameras with direct registration of MCP electrons with the Timepix ASIC metal pads have been also developed and tested [25, 26].
In applications which require single-photon detection, this sensitivity is achieved by coupling the camera to an image intensifier which employs a fast (typically P47 type) scintillator. This allows detection of individual optical photons with high spatial and temporal precision. As a result, time-stamping optical cameras, such as the Chronos Phoebe Timepix3CAM from ASI, have emerged as a versatile platform for investigating low-light, time-resolved, and photon-statistical phenomena in both fundamental research and applied science. The timing resolution is one of most important figures of merit for the time-stamping cameras. Here we describe the first tests of an intensified optical Timepix4 camera, focusing on its superior timing resolution compared to the previous version of the camera with Timepix3 chip [2].
There are several types of imaging cameras capable of single-photon detection, e.g., intensified time-stamping complementary metal-oxide-semiconductor cameras (iCMOS) [27], intensified or electron-multiplying charge-coupled device cameras (iCCD or EMCCD) [28, 29, 30, 31, 32, 33, 34], cameras based on direct registration of charge generated by micro-channel plates (MCP) in Timepix chips [35, 36] and resistive anodes (PhotonPixTM camera) [37], and cameras based on single-photon avalanche diodes (SPADs) [38, 39, 40, 41, 42, 43, 44], and we refer the readers to existing reviews of the subject [45, 46].
2 Methods
For the tests described here, we developed an optical camera based on the Timepix4 chip bump-bonded to a silicon sensor that is adapted for optical photons. This sensor is identical to the one used in optical cameras based on Timepix3. We developed a housing with a C-mount flange, so that an intensifier or lenses could be correctly mounted and focused.
The testing was performed with a pulsed blue 450 nm laser that illuminated a small area of the optical Timepix4 camera field of view. This was done both for intensified and non-intensified camera configurations. The timing studies relied on measurements of a delay between the laser pulse and the time-stamped pixel response. The experimental setup is schematically shown in the left part of Figure 1, and we describe it and its essential components below.
2.1 Timepix4 readout chip
The Timepix4 is a next-generation pixel readout chip developed by the Medipix4 collaboration, designed for high-resolution, time-resolved particle and X-ray detection [47]. It features a matrix of 512 × 448 pixels with a 55 m × 55 m pixel pitch, and offers simultaneous measurement of time-of-arrival (ToA) and time-over-threshold (ToT) per pixel. The chip supports data-driven readout, enabling each pixel to independently report events without the need for a global trigger, which is critical for sparse, high-rate environments. Timepix4 achieves timing resolution better than 200 ps when used with fast sensors, and supports multi-hit capability [48]. Its back-end architecture supports hit rates exceeding 180 Mhits/cm²/s, making Timepix4 well suited for high-throughput, time-resolved imaging. Timepix4 measures time with granularity of 195 ps corresponding to a best-case RMS resolution of about 60 ps. For comparison, the Timepix3 time bin size is 1.56 ns, corresponding to a time resolution limit of 450 ps.
2.2 Optical Timepix cameras
Optical cameras based on the Timepix3 chip have become a powerful tool for time-resolved photon detection in the visible and near-infrared domains used in a multitude of applications. The Timepix3 chip features a data-driven readout architecture, enabling each of its 256 × 256 pixels (with 55 µm pitch) to independently record both the time-of-arrival (ToA) and time-over-threshold (ToT) of incoming signals with a timing resolution of 1.56 ns and some energy discrimination via pulse-height ToT information [49].
In optical imaging applications, these cameras are typically coupled to optical sensors with high quantum efficiency in the 400-950 nm wavelength range [50]. The optical data-driven time-stamping concept was originally developed in 2015 for ion imaging in VMI applications when the first optical sensors compatible with the Timepix2 chip had been developed and produced [1]. For single-photon applications, image intensifiers are employed, which convert single photons into fast light flashes that can be detected by the optical sensor with high efficiency. This configuration enables single-photon sensitivity, nanosecond timing, and spatial resolution within the same platform. Since the intensifier and camera are completely mechanically decoupled, the same camera can be used with different intensifiers, which makes this configuration particularly flexible.
The first single-photon-sensitive Timepix camera was used for lifetime imaging in 2017 [8] and the Timepix3 camera was first tested for quantum applications in 2019 [15]. At present, optical Timepix3 cameras are increasingly used in fields such as quantum optics, ion and electron imaging, fluorescence lifetime imaging, and time-resolved spectroscopy, where precise spatio-temporal characterization of low-light signals is essential. Their ability to operate in event-driven mode without the need for external triggering makes them particularly well suited for detecting sparse, asynchronous photon events with minimal dead time.
In the optical Timepix4 camera used for these experiments, a Timepix3 optical sensor, which is smaller in area than the full Timepix4 chip, was bump-bonded to Timepix4 and read out with SPIDR4 electronics. The leakage current of the sensor was larger than usual, which is likely caused by the edges of the sensor touching parts of the Timepix4 solder bump array. We therefore kept the bias voltage limited in the experiments. The camera had mounting provisions for the image intensifier packaged as Photonis CricketTM, which integrates the intensifier, power supply, and back-end optics between the intensifier output and optical sensor. The camera also implements up to four time-digital-converters (TDC) employing some of peripheral Timepix4 pixels for precise time-stamping of external signals. We employed it to time-stamp the pulse generator synchronization signal used to trigger the 90 ps laser flash.
2.3 Optical setup
The 450 nm laser (Thorlabs GSL45A) was triggered with a pulse generator at 190 kHz. The laser pulse had the minimal allowed pulse duration of 90 ps. The produced light flash had provisions for attenuation before collimation on to the camera. In practice, the attenuation was used only in configuration with the intensifier. The right half of Figure 1 shows the camera (x,y) occupancy map with the intensifier for a well focused beam.
3 Results
The primary goal of this study was to evaluate the achievable time resolution for the Timepix4 camera in optical configurations. The key question here would be to quantify the improvement of temporal resolution due to the improved time binning of the readout chip, from 1.56 ns in Timepix3 to 195 ps in Timepix4, so an improvement by a factor of eight. However, it is not clear at all if this improvement can be realized as the signal amplification in the intensifier involves light emission of the P47 scintillator, which may limit the achievable resolution improvement due to its relatively slow time response with risetime of 7 ns [51]. The below study is the first attempt to quantify this effect.
3.1 Non-intensified camera
First we performed tests of the optical Timepix4 camera without the intensifier. In this configuration the camera is not single-photon sensitive, but detection can be achieved by a sufficiently bright flash. After adjusting the focusing a small spot corresponding to several pixels was illuminated on the sensor. We selected a single pixel in this spot and all presented below results were obtained for this single pixel so we did not need to account for the pixel-to-pixel variability issues.
The blue light with the wavelength of 450 nm is absorbed near the surface in silicon with absorption depth of about 0.5 micron. The produced carriers (holes as the sensor is p-on-n type) need to drift through the full thickness of depleted sensor to the Timepix4 pixels. The drift time would depend on the applied bias voltage and possibly on the properties of the surface, which may require additional bias voltage to be fully depleted. Diffusion of the drifting holes may affect the time resolution at 100 ps level but it is not expected to contribute significantly to the time resolution.
Figure 2 shows the two-dimensional distribution of time differences between the laser synchronization signal and ToA of one of the hit Timepix4 pixels activated with a flash of light from the laser versus the pixel ToT. We see that the ToT of the used pixel is contained in the 3500-3800 ns range and so does not vary much. We note that this value is due to direct illumination of the pixel during a single 90 ps laser flash.
The left part of Figure 3 shows the time difference distribution for a pixel activated with a flash of light from the laser without any TOT selection while the right graph shows the same but with TOT selection in a 25 ns slice. The 25 ns ToT slice is indicated with red dashed lines in Figure 2. The time resolution (RMS) was estimated by fitting the Gaussian function to the time difference distribution. The resolution was equal to 318 ps for the whole dataset and to 272 ps for a selected 25 ns slice of ToT values. The measurements were performed at a bias voltage of 50 V.
3.2 Intensified camera
We used the intensifier from Photonis with hi-QE-blue photocathode, chevron (double-layer) MCP at its maximum gain of about and P47 scintillator. The intensifier was used in the CricketTM configuration with the laser flash directed to the intensifier photocathode. In this case the laser pulse had to be strongly attenuated to provide a reasonable rate of single photons. The typically activated area of several pixels is shown in the right part of Figure 1.
The P47 scintillator emits light over a spectrum with maximum at 430 nm and approximate range of 390-490 nm [51], and, thus, it behaves similarly to the used 450 nm laser (in terms of photon absorption) with photon conversions near the silicon sensor surface and considerable drift time through the full sensor depth. The left part of Figure 4 shows the two-dimensional distribution of time differences between the laser synchronization signal and ToA of one of Timepix4 pixels versus ToT. The measurements were performed at a bias voltage of 70 V. The distribution clearly shows the expected dependence on ToT of the delay due to the timewalk effect. The time resolution for ToT selection in the range of 2000-2200 ns was estimated from the fit, as shown in the right part of Figure 4, to be equal to 0.83 ns. This demonstrates that we can achieve sub-nanosecond resolution for the intensified (and so single-photon sensitive) optical Timepix4 camera if the intensifier gain is large enough.
The ToT distribution is covering a wide range from small to large ToT values due to substantial gain variations of the MCP response to single photons, which is characteristic of any intensifier. We note that the largest ToT values in the intensified case reach the ToT range of non-intensified case and, therefore, the time resolutions can be directly compared for these two cases of similar ToT values. We performed the fit and estimated the time resolution for the intensified case with ToT in the range [3650-3675] at 550 ps. This needs to be compared to 272 ps in the non-intensified case in the right part of Figure 3. The difference in time jitter between these two cases can be explained mostly by the P47 scintillator contribution; we reckon that other possible contributions from the photocathode-to-MCP transit and in MCP itself are considerably smaller. We consider this as a viable confirmation that there are no fundamental limitations for intensified optical Timepix4 cameras to deliver the time resolution around 500 ps with high enough optical gain.
3.3 Timewalk correction
The timewalk terminology is used to describe the dependence of the timing response on the signal amplitude. For linear discriminators such as used in Timepix4 pixels larger signals cross the threshold earlier than signals with smaller amplitude. This effect can be studied and corrected as previously demonstrated with both Timepix3 and Timepix4 chips [52, 53, 3].
We performed timewalk correction by fitting the time difference distributions with a Gaussian function in 25 ns slices of ToT. Figure 5 shows results of the fit for the mean values and sigmas as function of ToT, respectively, in the left and right parts of the figure. The measurements were performed at a bias voltage of 70 V. We note that for the highest values of ToT the sigma value of the Gaussian fit and therefore the timing resolution for these ToT values is around 500-600 ps, which is very good indeed, also meaning that other contributing to the resolution factors (surface charge collection, diffusion in silicon and scintillator response) are below this value as discussed earlier.
After correction, the time difference distribution was fit with a Gaussian function, selecting ToT larger than 500 ns, which accounted for about 82% of events. The fit is shown in the right part of Figure 6, resulting in a time resolution of 1.42 ns. The left part of Figure 6 shows the two-dimensional distribution of time difference versus ToT after the timewalk correction. Projection of this distribution on the horizontal axis was used to produce the histogram in the right part of the figure.
3.4 Dependence on bias voltage
Photons emitted by the P47 scintillator in the intensifier are converted to photoelectrons near the sensor surface. Charge collection from the surface depends on depletion of silicon in this area and, therefore, depends on the allied bias voltage. We measured the dependence of the collection time behaviour on the bias voltage by acquiring datasets for six different voltages: 20, 30, 40, 50, 60 and 70 V in the configuration with intensifier. Figure 7 shows the ToT versus time difference distributions and individual time difference distributions for all values of the bias voltage. We see that the measured time difference decreases from over 100 ns values for bias of 20 V when the sensor may not be fully depleted to the values around 50 ns for maximum bias voltage of 70 V. The latter delay includes the drift time through 300 micron of depleted silicon and additional time offset between the signals. In all cases the timewalk effect is clearly visible. The time difference distributions in the right part of Figure 7 were selected with ToT in the 2000 - 2200 ns range indicated with red lines in the left part of the figure.
The left and right panels of Figure 8 present, respectively, the time delay between the laser synchronization signal and the ToA of a single Timepix4 pixel activated by a laser flash, and the corresponding time resolution as a function of bias voltage, both obtained with a ToT selection in a 200 ns window from 2000 to 2200 ns. Assuming the variation in delay with applied voltage is fully caused by the changing drift velocity of the holes through the sensor, we can describe the dependence as
with the sensor thickness, the constant offset due to other effects and the hole mobility. This function is fitted to the delay in the left part of Figure 8, yielding a mobility of (3.7 0.2)102 cm2V-1s-1, which is in good agreement with literature values at temperatures slightly above room temperature [54].
The time resolution for the ToT selections, as shown in the right part of Figure 8, reaches 0.83 ns at 70 V and remains better than 1 ns for a bias voltage as low as 40 V. These results indicate that the ultimate timing performance of the intensified optical camera with Timepix4 readout can be probed at higher bias voltages, which enable prompt charge carrier collection, and at the highest achievable ToT values, which minimize the impact of timewalk corrections and their associated jitter.
An interesting observation is also that the time resolution exhibits a dependence on the applied bias voltage even for identical ToT values. This behavior suggests sensor-related effects, such as, for example, increased time jitter arising from charge diffusion during carrier drift in the depleted silicon region. It is worth noting that the presented distributions are pixel-based (produced prior to clustering as required by the timewalk correction procedure) and therefore do not fully represent the intrinsic single-photon time jitter. A more detailed investigation of these effects is left for future work.
4 Conclusions
Timepix3 and Timepix4 are both hybrid-pixel, event-driven readout chips that, when coupled to an optical sensor and image intensifier, enable single-photon, time-resolved optical imaging. Timepix3 provides a 256×256 array (55 µm pitch), 1.56 ns time-of-arrival (ToA), and per-pixel time-over-threshold (ToT). Timepix4 enlarges the matrix to 512×448 at the same pitch ( 3.5× area), improves ToA binning to 195 ps, and sustains considerably higher hit rates. Both output ToT, but Timepix4’s upgraded front-end yields more accurate amplitude estimates. Overall, Timepix4 offers superior timing, throughput, and scalability for high-flux, sub-nanosecond optical experiments.
Table 1 presents a comparison of timing resolutions (RMS) achieved in this study for different configurations of Timepix4 camera. The resolution deteriorates after addition of the intensifier mostly due to appearance of a scintillator as an intermediate step in the amplification chain. The P47 scintillator typically used in fast image intensifiers has a rise time of about 7 ns and, therefore, can limit the time resolution in situations with insufficient gain. Faster scintillators are becoming available [51, 55], which can be used in intensifiers and also for similar MCP-based amplification schemes employed for ion and electron detection.
| Configuration | Time resolution [ns] | Comments |
| direct flash: no ToT selection | 0.32 | ToT range 300 ns |
| direct flash: ToT selection | 0.27 | ToT range 25 ns |
| intensifier: mild ToT selection + timewalk correction | 1.42 | large detection efficiency |
| intensifier: high ToT selection | 0.55 | small detection efficiency |
In summary, we presented the first characterization of an optical camera based on the Timepix4 chip coupled to an optical silicon sensor and image intensifier, demonstrating its potential for sub-nanosecond scale, time-resolved single-photon imaging. The system achieved time resolutions down to about 0.3 ns without the intensifier and as low as 0.55 ns with intensifier in the single-photon regime and selection of signals with high amplitude. For a more inclusive case of amplitude selections covering 82% of events the resolution with intensifier is equal to 1.4 ns, which is still a factor of two to three better than the resolution achieved in the past with the Timepix3 cameras [3, 2]. These measurements validate the capability of the Timepix4-based optical cameras to perform fast, high-resolution optical time-stamping. The observed performance marks a significant improvement over previous Timepix3-based systems, particularly in terms of timing precision, pixel count and readout rate.
Our results also highlight key factors influencing timing performance, including the sensor bias voltage, intensifier scintillator properties, and timewalk effects. After applying a timewalk correction and optimizing selection criteria, we demonstrated sub-nanosecond timing with intensified readout - a performance level suitable for single-photon imaging in quantum optics, ultrafast imaging, and time-correlated photon counting experiments. With its high throughput and improved timing resolution, the Timepix4 camera is well positioned as a scalable platform for next-generation time-resolved optical measurements across a broad range of scientific domains.
Acknowledgements.
This research was supported by the Czech Science Foundation (GACR) under Project No. 25-15534M and Czech Ministry of Education, Youth and Sports Project No. LM2023040 CERN-CZ. We thank Lou-Ann Pestana De Sousa for help with preparation of manuscript.References
- [1] Fisher-Levine, M. and Nomerotski, A., “TimepixCam: a fast optical imager with time-stamping,” Journal of Instrumentation 11(03), C03016 (2016).
- [2] Nomerotski, A., Chekhlov, M., Dolzhenko, D., Glazenborg, R., Farella, B., Keach, M., Mahon, R., Orlov, D., and Svihra, P., “Intensified Tpx3Cam, a fast data-driven optical camera with nanosecond timing resolution for single photon detection in quantum applications,” Journal of Instrumentation 18, C01023 (Jan. 2023).
- [3] Zhao, A., van Beuzekom, M., Bouwens, B., Byelov, D., Chakaberia, I., Cheng, C., Maddox, E., Nomerotski, A., Svihra, P., Visser, J., Vrba, V., and Weinacht, T., “Coincidence velocity map imaging using tpx3cam, a time stamping optical camera with 1.5 ns timing resolution,” Review of Scientific Instruments 88 (Nov. 2017).
- [4] Albrechtsen, S. H., Schouder, C. A., Viñas Muñoz, A., Christensen, J. K., Engelbrecht Petersen, C., Pi, M., Barranco, M., and Stapelfeldt, H., “Observing the primary steps of ion solvation in helium droplets,” Nature 623, 319–323 (Nov. 2023).
- [5] Sandström, E., Huysmans, P., Giskes, F., Laeven, P., Van Nuffel, S., Heeren, R. M. A., and Anthony, I. G. M., “Improvements in fast mass microscopy for large-area samples,” Analytical Chemistry 96, 18037–18042 (Oct. 2024).
- [6] Bromberger, H., Passow, C., Pennicard, D., Boll, R., Correa, J., He, L., Johny, M., Papadopoulou, C. C., Tul-Noor, A., Wiese, J., Trippel, S., Erk, B., and Küpper, J., “Shot-by-shot 250 khz 3d ion and mhz photoelectron imaging using timepix3,” Journal of Physics B: Atomic, Molecular and Optical Physics 55, 144001 (June 2022).
- [7] Roberts, A., Svihra, P., Al-Refaie, A., Graafsma, H., Küpper, J., Majumdar, K., Mavrokoridis, K., Nomerotski, A., Pennicard, D., Philippou, B., Trippel, S., Touramanis, C., and Vann, J., “First demonstration of 3d optical readout of a TPC using a single photon sensitive timepix3 based camera,” Journal of Instrumentation 14, P06001–P06001 (June 2019).
- [8] Hirvonen, L. M., Fisher-Levine, M., Suhling, K., and Nomerotski, A., “Photon counting phosphorescence lifetime imaging with timepixcam,” Review of Scientific Instruments 88(1), 013104 (2017).
- [9] Sen, R., Hirvonen, L. M., Zhdanov, A., Svihra, P., Andersson-Engels, S., Nomerotski, A., and Papkovsky, D., “New luminescence lifetime macro-imager based on a tpx3cam optical camera,” Biomed. Opt. Express 11, 77–88 (Jan 2020).
- [10] Sen, R., Zhdanov, A. V., Bastiaanssen, T. F. S., Hirvonen, L. M., Svihra, P., Fitzgerald, P., Cryan, J. F., Andersson-Engels, S., Nomerotski, A., and Papkovsky, D. B., “Mapping o2 concentration in ex-vivo tissue samples on a fast PLIM macro-imager,” Scientific Reports 10 (Nov. 2020).
- [11] D'Amen, G., Keach, M., Nomerotski, A., Svihra, P., and Tricoli, A., “Novel imaging technique for -particles using a fast optical camera,” Journal of Instrumentation 16, P02006–P02006 (Feb. 2021).
- [12] Losko, A. S., Han, Y., Schillinger, B., Tartaglione, A., Morgano, M., Strobl, M., Long, J., Tremsin, A. S., and Schulz, M., “New perspectives for neutron imaging through advanced event-mode data acquisition,” Scientific Reports 11(1) (2021).
- [13] Yang, J., Zhou, J., Jiang, X., Tan, J., Zhang, L., Zhou, J., Zhou, X., Yang, W., Xia, Y., Chen, J., Sun, X., Zhang, Q., Li, J., Sun, Z., and Chen, Y., “A novel energy resolved neutron imaging detector based on a time stamping optical camera for the CSNS,” Nuclear Instruments and Methods in Physics Research Section A: Accelerators, Spectrometers, Detectors and Associated Equipment 1000, 165222 (June 2021).
- [14] Wolfertz, A., Gustschin, A., Schulz, M., Long, A. M., Khaplanov, A., Hirsh, T. Y., Nomerotski, A., Morgano, M., Tremsin, A., Mauri, G., Sykora, G. J., and Losko, A., “Lumacam: a novel class of position-sensitive event mode particle detectors using scintillator screens,” Scientific Reports 14 (Dec. 2024).
- [15] Ianzano, C., Svihra, P., Flament, M., Hardy, A., Cui, G., Nomerotski, A., and Figueroa, E., “Fast camera spatial characterization of photonic polarization entanglement,” Scientific Reports 10 (Apr 2020).
- [16] Zhang, Y., England, D., Nomerotski, A., Svihra, P., Ferrante, S., Hockett, P., and Sussman, B., “Multidimensional quantum-enhanced target detection via spectrotemporal-correlation measurements,” Phys. Rev. A 101, 053808 (May 2020).
- [17] Svihra, P., Zhang, Y., Hockett, P., Ferrante, S., Sussman, B., England, D., and Nomerotski, A., “Multivariate discrimination in quantum target detection,” arXiv:2005.00612 (May 2020).
- [18] Nomerotski, A., Keach, M., Stankus, P., Svihra, P., and Vintskevich, S., “Counting of Hong-Ou-Mandel Bunched Optical Photons Using a Fast Pixel Camera,” Sensors 20(12) (2020).
- [19] Zhang, Y., England, D., Nomerotski, A., and Sussman, B., “High speed imaging of spectral-temporal correlations in hong-ou-mandel interference,” Optics Express 29, 28217 (Aug. 2021).
- [20] Gao, X., Zhang, Y., D’Errico, A., Heshami, K., and Karimi, E., “High-speed imaging of spatiotemporal correlations in hong-ou-mandel interference,” Optics Express 30, 19456 (May 2022).
- [21] Zhang, Y., Orth, A., England, D., and Sussman, B., “Ray tracing with quantum correlated photons to image a three-dimensional scene,” Physical Review A 105 (Jan. 2022).
- [22] Zhukas, L. A., Svihra, P., Nomerotski, A., and Blinov, B. B., “High-fidelity simultaneous detection of a trapped-ion qubit register,” Physical Review A 103 (June 2021).
- [23] Zhukas, L. A., Millican, M. J., Svihra, P., Nomerotski, A., and Blinov, B. B., “Direct observation of ion micromotion in a linear paul trap,” Physical Review A 103 (Feb. 2021).
- [24] Kato, A., Goel, A., Lee, R., Ye, Z., Karki, S., Liu, J. J., Nomerotski, A., and Blinov, B. B., “Two-tone doppler cooling of radial two-dimensional crystals in a radio-frequency ion trap,” Physical Review A 105 (Feb. 2022).
- [25] Vallerga, J., Tremsin, A., DeFazio, J., Michel, T., Alozy, J., Tick, T., and Campbell, M., “Optical mcp image tube with a quad timepix readout: initial performance characterization,” Journal of Instrumentation 9(05), C05055 (2014).
- [26] Bolzonella, R., Alozy, J., Ballabriga, R., Biesuz, N., Campbell, M., Cavallini, V., Ramusino, A. C., Fiorini, M., Franzoso, E., Guarise, M., Cudie, X. L., Romolini, G., and Saputi, A., “Development and characterization of hybrid mcp-pmt with embedded timepix4 asic used as pixelated anode,” Nuclear Instruments and Methods in Physics Research Section A: Accelerators, Spectrometers, Detectors and Associated Equipment 1082, 170965 (Feb. 2026).
- [27] Jachura, M. and Chrapkiewicz, R., “Shot-by-shot imaging of Hong–Ou–Mandel interference with an intensified sCMOS camera,” Optics letters 40(7), 1540–1543 (2015).
- [28] Jost, B. M., Sergienko, A. V., Abouraddy, A. F., Saleh, B. E., and Teich, M. C., “Spatial correlations of spontaneously down-converted photon pairs detected with a single-photon-sensitive CCD camera,” Optics Express 3(2), 81–88 (1998).
- [29] Brida, G., Caspani, L., Gatti, A., Genovese, M., Meda, A., and Berchera, I. R., “Measurement of Sub-Shot-Noise Spatial Correlations without Background Subtraction,” Phys. Rev. Lett. 102, 213602 (May 2009).
- [30] Zhang, L., Neves, L., Lundeen, J. S., and Walmsley, I. A., “A characterization of the single-photon sensitivity of an electron multiplying charge-coupled device,” Journal of Physics B: Atomic, Molecular and Optical Physics 42(11), 114011 (2009).
- [31] Fickler, R., Krenn, M., Lapkiewicz, R., Ramelow, S., and Zeilinger, A., “Real-time imaging of quantum entanglement,” Scientific reports 3(1), 1914 (2013).
- [32] Avella, A., Ruo-Berchera, I., Degiovanni, I. P., Brida, G., and Genovese, M., “Absolute calibration of an EMCCD camera by quantum correlation, linking photon counting to the analog regime,” Optics letters 41(8), 1841–1844 (2016).
- [33] Reichert, M., Sun, X., and Fleischer, J. W., “Quality of spatial entanglement propagation,” Physical Review A 95(6), 063836 (2017).
- [34] Moreau, P.-A., Toninelli, E., Gregory, T., and Padgett, M. J., “Imaging with quantum states of light,” Nature Reviews Physics 1(6), 367–380 (2019).
- [35] Vallerga, J., Tremsin, A., DeFazio, J., Michel, T., Alozy, J., Tick, T., and Campbell, M., “Optical mcp image tube with a quad timepix readout: initial performance characterization,” Journal of Instrumentation 9, C05055–C05055 (May 2014).
- [36] Alozy, J., Biesuz, N., Campbell, M., Cavallini, V., Cotta Ramusino, A., Fiorini, M., Guarise, M., and Llopart Cudie, X., “Development of a single-photon imaging detector with pixelated anode and integrated digital read-out,” Journal of Instrumentation 17, C06007 (June 2022).
- [37] Karl, S., Leopold, V., Richter, S., Prokazov, Y., Turbin, E., Sintotskiy, G., Orlov, D., and von Zanthier, J., “High-throughput single photon detection for effective stellar intensity interferometry,” in Advanced Photon Counting Techniques XIX , Itzler, M. A., McIntosh, K. A., and Bienfang, J. C., eds., 2, SPIE (May 2025).
- [38] Charbon, E., “Single-photon imaging in complementary metal oxide semiconductor processes,” Philosophical Transactions of the Royal Society A: Mathematical, Physical and Engineering Sciences 372(2012), 20130100 (2014).
- [39] Perenzoni, M., Pancheri, L., and Stoppa, D., “Compact SPAD-based pixel architectures for time-resolved image sensors,” Sensors 16(5), 745 (2016).
- [40] Gasparini, L., Bessire, B., Unternährer, M., Stefanov, A., Boiko, D., Perenzoni, M., and Stoppa, D., “SUPERTWIN: towards 100kpixel CMOS quantum image sensors for quantum optics applications,” in Quantum Sensing and Nano Electronics and Photonics XIV , 10111, 404–414, SPIE (2017).
- [41] Bruschini, C., Homulle, H., Antolovic, I. M., Burri, S., and Charbon, E., “Single-photon avalanche diode imagers in biophotonics: review and outlook,” Light: Science & Applications 8(1), 87 (2019).
- [42] Morimoto, K., Ardelean, A., Wu, M.-L., Ulku, A. C., Antolovic, I. M., Bruschini, C., and Charbon, E., “Megapixel time-gated SPAD image sensor for 2D and 3D imaging applications,” Optica 7(4), 346–354 (2020).
- [43] Lubin, G., Tenne, R., Ulku, A. C., Antolovic, I. M., Burri, S., Karg, S., Yallapragada, V. J., Bruschini, C., Charbon, E., and Oron, D., “Heralded spectroscopy reveals exciton–exciton correlations in single colloidal quantum dots,” Nano letters 21(16), 6756–6763 (2021).
- [44] Wojtkiewicz, M., Rae, B., and Henderson, R. K., “Review of Back-Side Illuminated 3-D-Stacked SPADs for Time-of-Flight and Single-Photon Imaging,” IEEE Transactions on Electron Devices (2024).
- [45] Hadfield, R. H., “Single-photon detectors for optical quantum information applications,” Nature Photonics 3, 696–705 (Dec. 2009).
- [46] Seitz, P. and Theuwissen, A. J., Single-photon imaging , vol. 160, Springer Science & Business Media (2011).
- [47] Llopart, X., Alozy, J., Ballabriga, R., Campbell, M., Casanova, R., Gromov, V., Heijne, E., Poikela, T., Santin, E., Sriskaran, V., Tlustos, L., and Vitkovskiy, A., “Timepix4, a large area pixel detector readout chip which can be tiled on 4 sides providing sub-200 ps timestamp binning,” Journal of Instrumentation 17, C01044 (Jan. 2022).
- [48] Heijhoff, K., Akiba, K., Ballabriga, R., van Beuzekom, M., Campbell, M., Colijn, A., Fransen, M., Geertsema, R., Gromov, V., and Llopart Cudie, X., “Timing performance of the timepix4 front-end,” Journal of Instrumentation 17, P07006 (July 2022).
- [49] Poikela, T., Plosila, J., Westerlund, T., Campbell, M., Gaspari, M. D., Llopart, X., Gromov, V., Kluit, R., Beuzekom, M. v., Zappon, F., Zivkovic, V., Brezina, C., Desch, K., Fu, Y., and Kruth, A., “Timepix3: a 65k channel hybrid pixel readout chip with simultaneous toa/tot and sparse readout,” Journal of Instrumentation 9, C05013–C05013 (May 2014).
- [50] Nomerotski, A., Chakaberia, I., Fisher-Levine, M., Janoska, Z., Takacs, P., and Tsang, T., “Characterization of timepixcam, a fast imager for the time-stamping of optical photons,” Journal of instrumentation 12(01), C01017 (2017).
- [51] Winter, B., King, S. J., Brouard, M., and Vallance, C., “A fast microchannel plate-scintillator detector for velocity map imaging and imaging mass spectrometry,” Review of Scientific Instruments 85 (Feb. 2014).
- [52] Tsigaridas, S., Beuzekom, M., Graaf, H., Hartjes, F., Heijhoff, K., Hessey, N., de Jong, P., and Prodanovic, V., “Timewalk correction for the timepix3 chip obtained with real particle data,” Nuclear Instruments and Methods in Physics Research Section A: Accelerators, Spectrometers, Detectors and Associated Equipment 930, 185–190 (June 2019).
- [53] Bolzonella, R., Alozy, J., Ballabriga, R., van Beuzekom, M., Biesuz, N., Campbell, M., Cardarelli, P., Cavallini, V., Coco, V., Cotta Ramusino, A., Fiorini, M., Gromov, V., Guarise, M., Llopart Cudie, X., Okamura, S., Romolini, G., Saputi, A., and Vitkovskiy, A., “Timing resolution performance of timepix4 bump-bonded assemblies,” Journal of Instrumentation 19, P07021 (July 2024).
- [54] Dorkel, J. and Leturcq, P., “Carrier mobilities in silicon semi-empirically related to temperature, doping and injection level,” Solid-State Electronics 24, 821–825 (Sept. 1981).
- [55] Zapadlík, O., Nikl, M., Polák, J., Průša, P., and Linhart, V., “Engineering of yag:ce to improve its scintillation properties,” Optical Materials: X 15, 100165 (Aug. 2022).